Write an equation of the straight line that is perpendicular to the line y=-5x + 8 and has x-intercept equal to 12.
Answer by ewatrrr(24785)


Hi


Write an equation of the straight line that is ⊥ to the line y=-5x + 8
y = (1/5)x + b | ⊥ Lines have slopes that are negative reciprocals of one another, that is -5*(1/5) = -1
y = (1/5)x + b |using order pair Pt(12,0) to solve for
0 = 12/5 + b
-12/5 = b
y = (1/5)x -12/5 OR x-5y = 12 in standard form
Note m = 1/5: as y increases/decreases by 1, x increases/decreases by 5